The production of heterotrophic flagellates, which are purely herbivorous, in an upwelling region... The production of heterotrophic flagellates, which are purely herbivorous, in an upwelling region is 0.35 gCm-2d-1, and the primary production is 1.5 gcm-2 d-1. What percentage of the primary production is consumed by the heterotrophic flagellates?
Answer #1

Hey, Welcome to Chegg. I hope my answer will be of some help to you. Thank you!

Primary production is 1.5 units

Production in upwelling region is 0.35 units

Part of primary production consumed = primary production- production in upwelling region

= 1.5-0.35 = 1.15 units

Percentage of primary production consumed = (consumed part ÷ total primary production) X 100

= (1.15 ÷ 1.5) X 100 = 0.76 x 100 = 76%